OpenStudy (anonymous):

42.5% of what is 25.16??

OpenStudy (ambi):

\(\Large\frac{42.5}{100} \times x = 25.16\) Solve for x

OpenStudy (anonymous):

1258/21

OpenStudy (anonymous):

@ambi

OpenStudy (ambi):

Not exactly correct, but it's very close

OpenStudy (anonymous):

Can you help me please

OpenStudy (anonymous):

@ambi

OpenStudy (ambi):

\(\Large x = \frac{25.16\times 100}{42.5}\)

OpenStudy (baru):

hi

OpenStudy (anonymous):

59.1/

OpenStudy (anonymous):

59.2

OpenStudy (baru):

yep

OpenStudy (ambi):

correct :D